OpenStudy (anonymous):

help please i will give medals!! i will also become fan!!!! If chelsea is trying to find the equation of a quadratic that has a focus of (1,3) and a directrix of y=-3, Describe to chelsea your preferred method for deriving the equation. make sure you use chelseas situation as a model to help her understand

OpenStudy (solomonzelman):

I forgot, guys help out please!

OpenStudy (anonymous):

It is really hard

OpenStudy (ranga):

Let (x,y) be a point on the parabola. The distance between (x, y) and focus = the distance between the point and the directrix.

OpenStudy (anonymous):

so what would that be I need the steps because that will just confuse me more!

OpenStudy (anonymous):

Distance is \[\sqrt{(y_2 - y_1)^2 + (x_2 - x_1)^2}\] So for example the distance between the parabola and the focus is \[\sqrt{(y-3)^2 + (x-1)^2}\]And the distance between the parabola and the directrix is \[\sqrt{(y+3)^2 + 0^2}\](since it's just a line in terms of y, you can just throw out x) Set those equal and rewrite it in standard form.

OpenStudy (anonymous):

thank you